# Graph Neural Networks for Next-Generation-IoT: Recent Advances and Open Challenges

### v2

@@ III Applications of GNN for NG Technologies in IoT Networks
[67] (2023) Massive MIMO Power control Node GNN
RIS [68] (2023) RIS aided communication Channel estimation Sub-graph GNN
@@
[67] (2023) RIS-enhanced downlink LEO satellite IoT networks Power control Node Unfolding GNN
Satellite [71] (2023) LEO satellite networks Satellite routing and network traffic optimization Node RL + MPNN
@@
[74] (2023) LEO satellite constellation Topology optimization Node RL + GNN
@@
[79] (2023) Integrated communication and sensing for vehicle communication Operation mode selection Node GNN
MEC [80] (2021) MEC support IoT networks Task off-loading Edge RL + GNN
@@
[82] (2022) MEC based UAV Task off-loading Node RL + GNN
@@
URLLC [89] (2023) Cellular network QoS violation probability minimization Node GNN
@@
[91] (2021) A factory automation scenario Packet loss probability minimization Node GNN
[92] (2024) massive URLLC (mURLLC) network Decoding error probability minimization Node GNN
Blockchain [158] (2023) A secure smart blockchain IoT network User privacy and data processing time Node GNN
@@
[95] (2020) IoT smart blockchain for healthcare applications Malicious node detection Node GNN
@@
[97] (2023) Federated learning Anomaly detection Node GNN
[98] 2024 Federated learning for UAV assisted MEC systems Classification inference Node GNN
[99] (2025) Federated learning for classification tasks Node classification tasks Node GNN
[100] (2025) Federated learning for edge computing Node classification tasks Node RL + GNN
[102] 2025 Decentralized tasks in wireless networks Channel impairments in GNNs Node GNN
@@
[104] 2024 Distributed fog servers for smart IoT services Communication efficiency Node GNN
@@
[38] (2023) RIS support federated learning Resource allocation Node GNN
[106] (2023) Federated learning over wireless network Power control Node GNN
[92] (2024) Distributed URLLC system Beamforming and power control Node GNN
@@ III-1 Background:
[…]
@@ III-1 Backgrounds:
As NG-IoT environments consist of diverse, distributed, and often resource-constrained edge devices, centralized data collection and processing pose critical challenges in terms of data privacy, communication costs, and model scalability. To address these issues, distributed paradigms, such as federated learning (FL) and decentralized structures, have gained significant attention. FL allows multiple edge devices to collaboratively train a shared model without exchanging raw data, preserving privacy and reducing communication overhead [182, 183, 184, 185, 186, 187]. Meanwhile, decentralized architectures eliminate the need for a central server by enabling peer-to-peer collaboration among devices [188, 63]. The integration of GNNs with distributed systems has emerged as a promising paradigm to meet the privacy, scalability, and efficiency requirements of next-generation IoT (NG-IoT) networks. Combining these two paradigms empowers edge nodes in NG-IoT systems to collaboratively learn graph-based models with minimal communication overhead and stronger privacy guarantees. This synergy opens two prominent research directions: (1) leveraging distributed architectures to optimize GNN(a) and (2) using GNNs to enhance distributed system in NG-IoT systems. Fig. 12 illustrates the integration of GNN with distributed frameworks in NG-IoT: (a) Federated learning-assisted GNN, where a central server aggregates local GNN models trained on distributed clients; (b) GNN-enabled decentralized system, where local agents make decisions collaboratively based on peer-to-peer graph reasoning without a central server.
@@ III-2 State-of-the-art:
In [101], Li et al. presented DHTGL, a distributed hierarchical temporal graph learning framework, reducing communication in IIoT systems. For decentralized training environments, Gao et al. [102] introduced AirGNNs, a communication-aware GNN architecture that incorporates wireless channel effects and enables over-the-air decentralized GNN training. The framework is robust to fading and noise and supports decentralized tasks such as multi-robot flocking and source localization. Li et al. [103] proposed a GCN-based dynamic topology selection strategy in vehicle-to-vehicle (V2V) networks to ensure reliable model sharing in decentralized federated learning (DFL), thereby improving training stability in mobile environments. In [104], Zeng et al. proposed FoGraph, a distributed GNN inference platform deployed over fog nodes. By leveraging compression-aware execution planning, FoGraph reduces communication bottlenecks compared to traditional cloud-based model serving and accelerates GNN inference by over 5×5$\times$5 ×.
@@ V-2 State-of-the-art:
Another promising scheme is constituted by the hybrid quantum graph neural network (HQGNN), which replaces the classical MLP in GNNs by quantum neural networks (QNNs) [123]. In details, as illustrated in Fig. 17, the HQGNN begins with graph input data, which is processed through hybrid graph convolutional layers that combine classical and quantum components. The quantum circuit performs updates on quantum states during training, effectively encoding node and edge relationships into the embedding space. By employing variational quantum circuits (VQC), QNNs are capable of facilitating quantum state updates during training, allowing for more efficient solution space exploration, while reducing the parameters and computational resources for large-scale graphs. This hybrid quantum-classical approach allows for more efficient exploration of the solution space than classical DNNs, hence reducing both the number of parameters and the computational resources required for processing large-scale graphs. By incorporating quantum circuits within the DQN, the computationally intensive parts of GNNs can be offloaded to quantum hardware, while the rest of the GNN relies on classical processors. This hybrid quantum-classical approach offloads computationally intensive tasks to quantum hardware, leaving the remaining calculations for classical processors, making it ideal for applications like NG networks. These include optimizing resource allocation, dynamic slicing, and traffic management. The simulated results presented in Fig. 8 highlighted the potential of the proposed hybrid quantum GNN. This approach demonstrates the potential of hybrid quantum GNNs and paves the way for future advances.

### v3

@@ III Applications of GNN for NG Technologies in IoT Networks
[70] (2023) Massive MIMO Power control Node GCN-WMMSE
RIS [71] (2023) RIS aided communication Channel estimation Sub-graph GCN + Transformer model
@@
Satellite [74] (2023) LEO satellite networks Satellite routing and network traffic optimization Node RL + MPGNN
@@
[77] (2023) LEO satellite constellation Topology optimization Node RL + MPGNN
@@
[82] (2023) Integrated communication and sensing for vehicle communication Operation mode selection Node HetGNN
MEC [83] (2021) MEC support IoT networks Task off-loading Edge RL + MPGNN
@@
[85] (2022) MEC based UAV Task off-loading Node RL + MPGNN
@@
URLLC [92] (2023) Cellular network QoS violation probability minimization Node REGNN
@@
[94] (2021) A factory automation scenario Packet loss probability minimization Node Edge-wise gate update + FNN
[95] (2024) Distributed URLLC system Decoding error probability minimization Node GraphSAGE
Blockchain [168] (2023) A secure smart blockchain IoT network User privacy and data processing time Node MPGNN
@@
[98] (2020) IoT smart blockchain for healthcare applications Malicious node detection Node GCN
@@
[100] (2023) Federated learning Anomaly detection Node MPGNN
[101] 2024 Federated learning for UAV assisted MEC systems Classification inference Node Simplified graph convolutional network
[102] (2025) Federated learning for classification tasks Node classification tasks Node GCN
[103] (2025) Federated learning for edge computing Node classification tasks Node RL + MPGNN
[105] 2025 Decentralized tasks in wireless networks Channel impairments in GNNs Node AirGNN
@@
[107] 2024 Distributed fog servers for smart IoT services Communication efficiency Node GCN, GAT, GraphSAGE
@@
[41] (2023) RIS support federated learning Resource allocation Node HetGNN
[109] (2023) Federated learning over wireless network Power control Node GCN
[95] (2024) Distributed URLLC system Beamforming and power control Node GraphSAGE
@@ III-1 Background:
Massive MIMO, RIS, satellite, THz, mobile edge computing, and URLLC systems are key NG technologies aimed at enhancing system performance, including energy, spectrum, and computational efficiency, as well as coverage. These technologies and their applications are comprehensively illustrated in Fig. 10, which highlights their roles in NG IoT networks. For instance, massive MIMO systems improve spectral efficiency by employing a large number of antennas, which allows for improved spatial multiplexing and user throughput [169, 170, 171, 172]. RIS creates programmable propagation environments capable of dynamically adjusting the phase, amplitude, and polarization of radio waves. This enhances coverage by reducing interference without relying on power amplifiers or complex signal processing [173, 174, 175, 176, 177]. Satellite networks have become crucial in supporting non-terrestrial networks by extending coverage to remote areas, hence providing global connectivity [178, 179, 180]. THz communication leverages the unlicensed spectrum ranging from 0.1 to $0.1\text${ to }10 THz for opening up new frequency bands. The abundance of free spectrum in this band circumvents spectrum scarcity, hence offering the high throughput needed for advanced applications like virtual reality and ultra-high-definition video [181, 182, 183, 184], albeit at a high path-loss and limited coverage. Mobile Edge Computing (MEC) brings data processing capabilities closer to end users by carrying out computation at the network edge instead of centrally, thus reducing latency and improving response times for time-sensitive applications. Harnessing MEC is essential in scenarios where quick decision-making and real-time analytics are needed, such as smart cities, autonomous vehicles, and industrial IoT environments [83, 84, 107, 85]. It helps offload computational tasks from centralized cloud servers, optimizing resource utilization and enhancing the overall network efficiency. Meanwhile, URLLC (Ultra-Reliable Low-Latency Communication) ensures low-latency, high-reliability communication for mission-critical applications like autonomous driving, robotics, and healthcare. It is crucial in near real-time scenarios, such as industrial automation and remote medical procedures, using bespoke resource allocation for meeting stringent performance requirements [185, 186]. Despite these advantages, the above sophisticated technologies face common challenges in managing high-dimensional spatial data and optimizing spectrum usage. GNNs are becoming popular in addressing these challenges by learning spatial dependencies, optimizing resource management, and maintaining efficient, low-latency communication in dynamic environments.
@@ III-1 Backgrounds:
As NG-IoT environments consist of diverse, distributed, and often resource-constrained edge devices, centralized data collection and processing pose critical challenges in terms of data privacy, communication costs, and model scalability. To address these issues, distributed paradigms, such as federated learning (FL) and decentralized structures, have gained significant attention. FL allows multiple edge devices to collaboratively train a shared model without exchanging raw data, preserving privacy and reducing communication overhead [192, 193, 194, 195, 196, 197, 198]. Meanwhile, decentralized architectures eliminate the need for a central server by enabling peer-to-peer collaboration among devices [199, 66]. The integration of GNNs with distributed systems has emerged as a promising paradigm to meet the privacy, scalability, and efficiency requirements of next-generation IoT (NG-IoT) networks. Combining these two paradigms empowers edge nodes in NG-IoT systems to collaboratively learn graph-based models with minimal communication overhead and stronger privacy guarantees. This synergy opens two prominent research directions: (1) leveraging distributed architectures to optimize GNN(a) and (2) using GNNs to enhance distributed system in NG-IoT systems. Fig. 12 illustrates the integration of GNN with distributed frameworks in NG-IoT: (a) Federated learning-assisted GNN, where a central server aggregates local GNN models trained on distributed clients; (b) GNN-enabled decentralized system, where local agents make decisions collaboratively based on peer-to-peer graph reasoning without a central server.
@@ III-2 State-of-the-art:
In [104], Li et al. presented DHTGL, a distributed hierarchical temporal graph learning framework, reducing communication in IIoT systems. For decentralized training environments, Gao et al. [105] introduced AirGNNs, a communication-aware GNN architecture that incorporates wireless channel effects and enables over-the-air decentralized GNN training. The framework is robust to fading and noise and supports decentralized tasks such as multi-robot flocking and source localization. Li et al. [106] proposed a GCN-based dynamic topology selection strategy in vehicle-to-vehicle (V2V) networks to ensure reliable model sharing in decentralized federated learning (DFL), thereby improving training stability in mobile environments. In [107], Zeng et al. proposed FoGraph, a distributed GNN inference platform deployed over fog nodes. By leveraging compression-aware execution planning, FoGraph reduces communication bottlenecks compared to traditional cloud-based model serving and accelerates GNN inference by over $5\times$.
@@ V-2 State-of-the-art:
Another promising scheme is constituted by the hybrid quantum graph neural network (HQGNN), which replaces the classical MLP in GNNs by quantum neural networks (QNNs) [126, 214]. In details, as illustrated in Fig. 17, the HQGNN begins with graph input data, which is processed through hybrid graph convolutional layers that combine classical and quantum components. The quantum circuit performs updates on quantum states during training, effectively encoding node and edge relationships into the embedding space. By employing variational quantum circuits (VQC), QNNs are capable of facilitating quantum state updates during training, allowing for more efficient solution space exploration, while reducing the parameters and computational resources for large-scale graphs. This hybrid quantum-classical approach allows for more efficient exploration of the solution space than classical DNNs, hence reducing both the number of parameters and the computational resources required for processing large-scale graphs. By incorporating quantum circuits within the DQN, the computationally intensive parts of GNNs can be offloaded to quantum hardware, while the rest of the GNN relies on classical processors. This hybrid quantum-classical approach offloads computationally intensive tasks to quantum hardware, leaving the remaining calculations for classical processors, making it ideal for applications like NG networks. These include optimizing resource allocation, dynamic slicing, and traffic management. The simulated results presented in Fig. 8 highlighted the potential of the proposed hybrid quantum GNN. This approach demonstrates the potential of hybrid quantum GNNs and paves the way for future advances.
@@ VII-A Lessons Learned and Future Research Directions
  • Privacy preservation in GNN-based NG-IoT systems presents unique challenges distinct from adversarial robustness. The message-passing mechanisms in GNNs may unintentionally expose critical structural patterns, such as influential node connections, bridge links, or tightly-knit communities, making generic privacy preservation techniques like simple anonymization or random link concealment ineffective. Such indiscriminate defenses often degrade model utility while failing to protect high-impact, context-sensitive relationships that adversaries are most likely to target. Future research must further advance context-aware privacy frameworks that selectively protect sensitive elements of NG-IoT networks while maintaining GNN utility. A key future direction is developing topology-aware DP mechanisms for graphs, since traditional DP methods inject excessive noise due to structural correlations. Recent advances demonstrate decentralized DP for subgraph statistics, personalized DP-GNNs enabling heterogeneous privacy budgets, and privacy-preserving embeddings [147, 146, 145]. These approaches collectively pave the way for context-aware privacy solutions that balance strong privacy guarantees with high utility in NG-IoT networks. Another direction is topology-aware graph perturbation, where measures like betweenness centrality or edge criticality guide selective link obfuscation to protect high-impact connections, while avoiding unnecessary anonymization [222]. Complementary to DP, homomorphic encryption [223, 224] enables privacy-preserving GCN inference by processing encrypted graph data on cloud servers. These methods protect both graph features and topology without revealing raw data, albeit at the cost of additional computational complexity.
@@
  • Semantic communications represent an emerging frontier for GNN integration, where learning-based models extract, process, and transmit meaning rather than raw data. Future research should explore how GNNs can support knowledge graph reasoning, semantic-aware encoding, and context-driven signal processing in NG-IoT applications. Papers by Zhang et al. [225], Guo et al. [226], and Hello et al. [227] offer early evidence of this synergy, but more robust GNN architectures are needed to maintain semantic fidelity under dynamic and resource-constrained conditions.